OpenStudy (anonymous):

what is the slope of the line that is perpendicular to the line that passes through the points (7,6) and (3,-4)

OpenStudy (anonymous):

first you want to find the slope of those points

OpenStudy (anonymous):

so how would i do that?

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay I got 2.5

OpenStudy (anonymous):

next you put it into point slope form or (y-y1)=m(x-x1)

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

well you found your slope m and pick one of your sets of points and plug them in for y1 and x1

OpenStudy (anonymous):

okay i got -2/5 for an answer, is that correct?

OpenStudy (anonymous):

y-6=2.5(x-7)

OpenStudy (tamtoan):

no, @katlin95, slope of the perpendicular line to a line with slope m is -1/m ...leave your slope in fraction form and see what -1/m = ?

OpenStudy (anonymous):

hhhm this is hard for me lol

OpenStudy (anonymous):

y-6=(1/2.5)x-((1/2.5)*7)

OpenStudy (tamtoan):

@katlin95, once you get slope m which you already found, just find what -1/m is :) that isn't hard is it?

OpenStudy (tamtoan):

the question ask for slope of perpendicular line, nickr is finding the equation, not the slope

OpenStudy (anonymous):

oh all you need is the slope my bad, didnt see that

OpenStudy (anonymous):

nice catch

OpenStudy (tamtoan):

@nickr np :) enjoy the night

OpenStudy (anonymous):

thanks same to you

OpenStudy (anonymous):

-2??? right?

OpenStudy (anonymous):

no your slope is 2.5 so your answer would be 1/2.5

OpenStudy (tamtoan):

your slope m is 10/4 = 5/2 ...so -1/m = -1/(5/2) = -2//5 don't write slope in decimal ...leave it as fraction, will make it much easier for you

OpenStudy (tamtoan):

and thanks nick :)

OpenStudy (anonymous):

i dont have that as an answer all i got was -2/5 for answer, which i said long before as my first answer

OpenStudy (anonymous):

thanks so much guys!

OpenStudy (anonymous):

yes you are correct trying to type to fast and forgot negative

OpenStudy (anonymous):

np sorry i messed up

OpenStudy (anonymous):

we all make mistake:)

OpenStudy (tamtoan):

nickr, don't you worry, i made mistakes all the time :)

OpenStudy (anonymous):

what now you dont lol